Use a right triangle where A =40 (degrees) and a = 4. solve the triangle, and round to 2 decimal places.
Thank You

Answer by swincher4391(1107)

Since there are 180º in a triangle, and we are given that it is a right triangle with angle 'a' being 40, that means that angle 'b' must be (90-40 = 50)º. We can then use the law of sines.
4/sin(40) = b/sin(50)
4sin(50) = bsin(40)
4sin(50)/sin(40) = b
b = 4.77
a^2 + b^2 = c^2
4^2 + (4.77)^2 = c^2
c = 6.22
Hope this helps!